# Quantifying Geospatial in the Common Crawl Corpus

### v1

@@ Abstract.
Large language models (LLMs) exhibit emerging geospatial capabilities, stemming from their pre-training on vast unlabelled text datasets that are often derived from the Common Crawl corpus. However, the geospatial content within CC remains largely unexplored, impacting our understanding of LLMs’ spatial reasoning. This paper investigates the prevalence of geospatial data in recent Common Crawl releases using Gemini, a powerful language model. By analyzing a sample of documents and manually revising the results, we estimate that between 1 in 5 and 1 in 6 documents contain geospatial information such as coordinates and street addresses. Our findings provide quantitative insights into the nature and extent of geospatial data within Common Crawl, and web crawl data in general. Furthermore, we formulate questions to guide future investigations into the geospatial content of available web crawl datasets and its influence on LLMs.
@@ 1. Introduction
Recent studies have found that large language models (LLMs) may have some knowledge and understanding of physical space. Gurnee and Tegmark (Gurnee and Tegmark, 2024) demonstrated that Llama-2 learned linear representations of space at various scales, such as landmarks and cities; the authors even identified specific nodes in the model’s architecture that are responsible for this behaviour. Bhandari et al. (Bhandari et al., 2023) probed Llama and several smaller language models for geocoding abilities and geospatial awareness, showing that larger and more sophisticated models have better abilities to synthesise geospatial knowledge from pre-training texts. Mai et al. (Mai et al., 2023) demonstrated superior performance of the GPT-family LLMs on toponym recognition and location description recognition tasks, with few-shot learning often performing on par with the specialised, state-of-the-art named entity recognition (NER) models.
@@
Given the importance of CC in the field of generative AI and the growing concerns over the biases and ethics of the generative AI models, it is not surprising the corpus has been analysed for undesirable content such as hate speech and sexually explicit content (Luccioni and Viviano, 2021; Raffel et al., 2020), and its derivative datasets were analysed for general quality (Kreutzer et al., 2022; Dodge et al., 2021; Birhane et al., 2021). While the geospatial bias of LLMs has been subject to multiple investigations (Moayeri and Feizi, 2024; Manvi et al., 2024; Fulman et al., 2024), to our knowledge, no attempts were made to look into the geospatial contents of CC.
In this paper, we aim to fill in this gap by quantifying the number of geospatial files found in recent CC releases, as well as estimating the share of traditional web documents (for example, HTML and XML files) that contain geospatial information, such as coordinates or street addresses. We believe that any debate about the emergent geospatial capabilities of LLMs cannot be complete without knowing the prevalence of geospatial content in the pre-training text corpus.
@@ 2. Related work
These works gave us insight into how researchers approach the problem of analysing and quantifying phenomena in large text corpora, and we propose our own method of quantifying geospatial in CC in the following section.
@@ 3. Method
Our aim is to estimate what share of documents in a typical CC release contains any geospatial information as defined in the next subsections. We deliberately avoid labelling and then quantifying the actual geospatial text as our delineation criteria and judgement of relevance would get extremely complex.
@@ 3.1. About Common Crawl
Common Crawl is a San-Francisco based non-profit organisation. Established in 2007, its goal is to ‘make wholesale extraction, transformation, and analysis of open web data accessible to researchers’ (Crawl, 2024). The released datasets comprise the world’s largest free corpus of scraped web data (Baack, 2024), estimated to contain about 250 billion pages, with between 3 and 5 billion added each month (Crawl, 2024).
The datasets are released every one-two months, and are freely available via S3 and HTTP protocols111111https://data.commoncrawl.org. Each release is divided into 300 Web Archive (WARC) chunks (files) for easier access, and accompanied by the index with relevant metadata on each file, such as file’s original URL, MIME-type, and content length. One can obtain individual scraped web documents from the WARC chunks directly from the Common Crawl servers, as we showed in our recent work (Ilyankou et al., 2024):
@@
In the code snippet above, we use requests Python library121212https://docs.python-requests.org/en/latest/index.html to retrieve an individual scraped web file from a CC warc_file using the Range HTTP header, only reading the bytes of the relevant file by knowing its location within the warc_file, including the offset from the start of the warc_file, and its length in bytes.
@@ 3.2. What constitutes geospatial
The vast majority of the files found in CC are standard web pages, such as HTML and XML pages. In fact, the share of CC files that are exclusively geospatial, such as .geojson or .kml, is only roughly 0.003-0.004%. Table 1 shows all geospatial MIME types that we were able to identify from the index of the three relevant CC releases. Therefore, our aim is to estimate the prevalence of geospatial content in predominantly non-geospatial files.
@@
| Total geo files | 89,693 | 94,290 | 139,126 |
| All docs in release | 2914M | 2930M | 3107M |
| Geo files as % of all docs | 0.003 | 0.003 | 0.004 |
In the Manual of Geospatial Science and Technology, Bossler defines geospatial data as ‘features tied to the surface of the earth by coordinates, addresses, or other means’ (Bossler, 2010). Thus, we focus on two elements:
@@
We do not count derivative location systems such as what3words131313https://what3words.com or geohash141414https://geohash.org/ as valid for our purposes. (In real life, these are typically accompanied by traditional addresses or coordinates, so we are likely to still count the document.) We exclude simple mentions of landmarks, such as ‘The White House’, unless they are accompanied by an address (e.g., ‘The White House, 1600 Pennsylvania Avenue NW, Washington, DC’ or even simply ‘The White House, Washington DC’). This is to ensure we don’t count documents that mention landmarks in passing.
We do not discriminate against real and fictional addresses. For example, we count ‘123 Main Street¡br/¿New York, NY 10001’ as a valid address even though we are certain it is used as a filler. Similarly, we do not check if latitude-longitude pairs are valid as long as it is clear the pair of numbers represents physical coordinates. We do, however, exclude the coordinates of the Null Island (0,0) as those are often used for map initialisation and are not meaningful otherwise (Juhasz and Mooney, 2022), as well as astronomical coordinates of celestial bodies where possible.
@@ 3.3. Identifying coordinates and addresses
In general, identifying coordinates can be performed using regular expressions, as those are typically expressed as pairs of numbers. For example, latitudes and longitudes are typically expressed as precise numbers between and , respectively. However, any regex pattern is likely to produce false positive matches, such as scientific numbers () or unusually formatted phone numbers (e.g., ). In addition, we occasionally see coordinates expressed in minutes and seconds, such as ‘N 56° 59’ 55” E 009° 19’ 33,7”’, as well as those expressed using various national grid systems, such as positive integers representing easting and northing in the British National Grid161616https://britishnationalgrid.uk/. Adding more regular expressions to capture all these cases is likely to produce even more false positives in different languages, which will be laborious to classify manually.
@@
In addition to varied formats, many texts are in non-Latin scripts, such as Cyrillic, Arabic, Greek, and Chinese/Japanese/Korean characters. Popular open-source software packages such as pyap171717https://github.com/vladimarius/pyap or libpostal181818https://github.com/openvenues/libpostal usually only work with shorter strings and are typically limited to particular countries. In 2010, Chang and Li (Chang and Li, 2010) proposed MapMarker, a sequence labelling algorithm coupled with pattern mining, to perform postal addresses extraction from web texts, achieving the F-score of 91%. In 2013, Schmidt et al. (Schmidt et al., 2013) suggested a hybrid approach using patterns and data from gazetteers to extract addresses from web documents, showing high recall and precision. However, in their approach HTML tags are removed from the documents, which is not ideal given that geospatial information can be housed inside <meta> and other tags, as well as inside JavaScript code snippets and JSON files. In 2018, Efremova et al. (Efremova et al., 2018) introduced a geo-tagging framework that employs n-grams, sliding windows, and a Support Vector Machine classifier to extract addresses from CC documents, achieving precision and recall of 91% and 93% respectively. Named entity recognition (NER) tools such as spaCy191919https://spacy.io/ can also be trained to extract addresses, yet this would require a labelled training set.
Inspired by the needle-in-a-haystack evaluation task for LLMs202020This popular GitHub repository contains needle-in-a-haystack benchmark tests for LLMs: https://github.com/gkamradt/LLMTest_NeedleInAHaystack, in which long-context window LLMs are tasked to retrieve particular bits of information from a long document, such as an article or a book, we decided to employ a language model to retrieve both coordinates and addresses from web documents (if they exist), and then manually review the results. LLMs are excellent zero-shot learners and have demonstrated superior performance in named entity recognition and text classification tasks (Xie et al., 2023; Zhou et al., 2024; Wei et al., 2022; Team, 2024), so we believe they are better suited to identify the presence of geospatial in multilingual CC documents than traditional, language-and country-specific methods.
@@
Gemini 1.5 is one of the few LLMs currently capable of handling inputs of up to 1 million tokens, and has shown to perform well in the needle-in-a-haystack retrieval task (Team, 2024). The context window consideration is important given the web documents found in CC can be lengthy. Figure 1 shows the histogram of lengths of sampled documents that we used in our analysis. While the median and means lengths range from ca. 17k to 55k tokens between the releases, the length distribution is heavily skewed to the right, with some documents reaching up to 1M tokens.
@@
The prevalence of geospatial data in CC documents
In this paper, we use the Gemini 1.5 Flash version, which has the most generous free tier API usage allowance of 1,500 requests per day. At the time of running our experiment, the pay-as-you-go tier with up to 10,000 requests per day has been announced, but is not yet available212121According to https://ai.google.dev/pricing, the pay-as-you-go pricing for Gemini 1.5 Flash and Pro will become available on May 30, 2024.
After some instruction-tuning on sampled web documents from CC, we adapted the following query:
@@
Below is an example of a Gemini response to the query for the document that contains both a pair of coordinates and a street address:
@@ 3.5. Determining sample size
However, in our task we are unable to utilise fast algorithms such as keyword lookup, n-grams, or K-Means. We are limited by the number of API requests we can make; in addition, each query takes Gemini between several seconds to several minutes to be processed. We need to choose the smallest yet scientifically sound sample size, and so we decided to use Cochran’s formula (1) (Bartlett et al., 2001). The formula is used to calculate the ideal sample size that estimates population with a desired level of precision and a particular confidence interval. It is commonly used in social sciences to estimate the number of people to survey. We use the version of Cochran’s formula modified for very large (or infinite) population because a single CC release contains billions of web pages:
@@
3.6. Revising Gemini responses
We manually reviewed each Gemini response that is longer than 1 or 2 tokens (‘None’ or ‘None.’, respectively) in order to check that we are satisfied with the quality of the examples returned, as well as to distinguish whether the document contains an address, or coordinates, or both.
@@ 3.6. Revising Gemini responses
In order to check for false negatives, that is, the documents for which Gemini returned None that may contain geospatial data, we manually reviewed a random sample of 100 such documents. Using Google Translate222222https://translate.google.com when needed, we were unable to identify street addresses or coordinates in all cases.

### v2

@@ Abstract.
Large language models (LLMs) exhibit emerging geospatial capabilities, stemming from their pre-training on vast unlabelled text datasets that are often derived from the Common Crawl (CC) corpus. However, the geospatial content within CC remains largely unexplored, impacting our understanding of LLMs’ spatial reasoning. This paper investigates the prevalence of geospatial data in recent Common Crawl releases using Gemini 1.5, a powerful language model. By analyzing a sample of documents and manually revising the results, we estimate that 18.7% of web documents in CC contain geospatial information such as coordinates and addresses. We find little difference in prevalence between Enlgish- and non-English-language documents. Our findings provide quantitative insights into the nature and extent of geospatial data in CC, and lay the groundwork for future studies of geospatial biases of LLMs.
@@ 1. Introduction
Recent studies have found that large language models (LLMs) may have some knowledge and understanding of physical space. Gurnee and Tegmark (Gurnee and Tegmark, 2024) demonstrated that Llama-2 learned linear representations of space at various scales, such as landmarks and cities; the authors even identified specific nodes in the model’s architecture that are responsible for this behaviour. Bhandari et al. (Bhandari et al., 2023) probed Llama and several smaller language models for geocoding abilities and geospatial awareness, showing that larger and more sophisticated models have better abilities to synthesise geospatial knowledge from pre-training texts. Mai et al. (Mai et al., 2023) demonstrated superior performance of the GPT-family LLMs on toponym recognition and location description recognition tasks, with few-shot learning often performing on par with the specialised, state-of-the-art named-entity recognition (NER) models.
@@
Given the importance of CC in the field of generative AI and the growing concerns over the biases and ethics of the generative AI models, it is not surprising the corpus has been analysed for undesirable content such as hate speech and sexually explicit content (Luccioni and Viviano, 2021; Raffel et al., 2020), and its derivative datasets were analysed for general quality (Kreutzer et al., 2022; Dodge et al., 2021; Birhane et al., 2021). While the geospatial bias of LLMs has been subject to multiple investigations (Moayeri and Feizi, 2024; Manvi et al., 2024; Fulman et al., 2024), sometimes producing disturbing results (such as that the ‘error rates are 1.5 times higher for countries from Sub-Saharan Africa compared to North American countries’ (Moayeri and Feizi, 2024) and ‘the LLMs exhibit geographic bias across objective and subjective topics, particularly discriminating against areas with lower socioeconomic conditions’ (Manvi et al., 2024)), to our knowledge, no attempts were made to look into the geospatial contents of CC.
In this paper, we aim to fill in this gap by quantifying the number of geospatial files found in recent CC releases, as well as estimating the share of traditional web documents (for example, HTML and XML files) that contain geospatial information, such as coordinates or street addresses. We believe that any debate about the emergent geospatial capabilities of LLMs, as well as the geospatial biases they demonstrate, cannot be complete without knowing the prevalence of geospatial content in the pre-training text corpus.
@@ 2. Related work
These works gave us insight into how researchers approach the problem of analysing and quantifying phenomena in large text corpora. We propose our own method of quantifying geospatial in CC in the following section.
@@ 3. Method
Our aim is to estimate what share of documents in a typical CC release contains any geospatial information as defined in Subsection 3.2. We deliberately avoid labelling and then quantifying the actual geospatial text as our delineation criteria and judgement of relevance would get extremely complex.
@@ 3.1. About Common Crawl
Common Crawl is a San-Francisco based non-profit organisation. Established in 2007, its goal is to ‘make wholesale extraction, transformation, and analysis of open web data accessible to researchers’ (Crawl, 2024a). The released datasets comprise the world’s largest free, multi-lingual corpus of scraped web data (Baack, 2024b), estimated to contain about 250 billion pages (over 9.5 petabytes), with between 3 and 5 billion pages added each month (Crawl, 2024a). Due to the dataset’s scale, it is sometimes incorrectly described as containing ‘the entire web’. As Baack (Baack, 2024b) points out, it is not even a representative sample of the web.
New CC releases are published every one-two months, and are freely available via S3 and HTTP protocols111111https://data.commoncrawl.org. Each release is divided into 300 Web Archive (WARC) chunks for easier access, and accompanied by the index with relevant metadata on each web document, such as the web document’s original URL, MIME-type, and content length. One can obtain individual web documents from the WARC chunks directly from the Common Crawl servers, as we demonstrated in our recent paper (Ilyankou et al., 2024):
@@
In the code snippet above, we use requests Python library121212https://docs.python-requests.org/en/latest/index.html to retrieve an individual scraped web document from a CC warc_file using the Range HTTP header, only reading the bytes of the relevant web document by knowing its location within the warc_file, including the offset from the start of the warc_file, and the web document’s length in bytes.
@@ 3.2. What constitutes geospatial
The vast majority of the files found in CC are standard web pages, such as HTML and XML pages. In fact, the share of CC files that are exclusively geospatial, such as .geojson or .kml, is only roughly 0.003-0.004%. Table 1 shows all geospatial MIME types that we were able to identify from the index of the three studied CC releases. Therefore, our main goal is to estimate the prevalence of geospatial content in predominantly non-geospatial files.
@@
| Total geospatial files | 89,693 | 94,290 | 139,126 |
| All documents in release | 2914M | 2930M | 3107M |
| Geo. files as % of all docs. | 0.003 | 0.003 | 0.004 |
According to the ISO’s definition of geographic data (which is essentially the same as geospatial data), it must contain ‘implicit or explicit reference to a location relative to the Earth’ (ISO, 2012). In the Manual of Geospatial Science and Technology, Bossler defines geospatial data as ‘features tied to the surface of the earth by coordinates, addresses, or other means’ (Bossler, 2010). In this paper, we focus on the first two reference types:
@@
While we agree that there exist numerous other, lesser used but perfectly valid ways to express locations precise enough to pinpoint a building or a business, such as ‘the cafe at the foot of the Eiffel tower’ or descriptors that use local landmarks and unique characteristics to identify places within informal settlements (references which Bossler may have classed as other means), after many discussions, we decided not to include them in our calculations for two reasons. Firstly, the threshold for what to include would become fuzzy. Secondly, we would be reliant on the LLM’s perception of spatial reasoning and understanding of cultural contexts and navigation to identify and return these examples, inevitably introducing bias to our results.
We do not count derivative location systems such as what3words131313https://what3words.com or geohash141414https://geohash.org/ as valid for our purposes. (In real life, these are typically accompanied by traditional addresses or coordinates, so we are likely to still count the documents.) We exclude simple mentions of landmarks, such as ‘the White House’, unless they are accompanied by an address (e.g., ‘The White House, 1600 Pennsylvania Avenue NW, Washington, DC’ or even simply ‘The White House, Washington DC’). This is to ensure we don’t count documents that mention landmarks in passing, although we recognise it can be argued that LLMs may learn the physical world from such examples as well.
We do not discriminate against real and fictional addresses. For example, we count ‘123 Main Street¡br/¿New York, NY 10001’ as a valid address even though we are certain it is used as a filler. Similarly, we do not check if latitude-longitude pairs are valid as long as it is clear the pair of numbers represents physical coordinates. We do, however, exclude the coordinates of the Null Island (0,0) as those are often used for map initialisation and are not meaningful otherwise (Juhasz and Mooney, 2022). We also exclude astronomical coordinates of celestial bodies where possible.
@@ 3.3. Identifying coordinates and addresses
"spatialReference": {
"wkid": 27700
"x": 263779.76690000016,
"y": 235218.1986999996
In general, identifying coordinates can be performed using regular expressions, as those are typically expressed as pairs of numbers. For example, latitudes and longitudes are expressed as real numbers between and , respectively. However, any regex pattern is likely to produce false positive matches, such as scientific numbers () or unusually formatted phone numbers (e.g., ). In addition, we occasionally see coordinates expressed in minutes and seconds, such as ‘N 56° 59’ 55” E 009° 19’ 33,7”’, as well as those expressed using various national grid systems, such as positive integers representing easting and northing in the British National Grid161616https://britishnationalgrid.uk/. Adding more regular expressions to capture all these cases is likely to produce even more false positives in different languages, which will be laborious to classify manually.
@@
[…]
In addition to varied formats, many texts are in non-Latin scripts, such as Cyrillic, Arabic, Greek, and Chinese/Japanese/Korean characters, occasionally even Unicode-escaped (see final example above). Popular open-source software packages such as pyap171717https://github.com/vladimarius/pyap or libpostal181818https://github.com/openvenues/libpostal usually only work with shorter strings and are typically limited to particular countries. In 2010, Chang and Li (Chang and Li, 2010) proposed MapMarker, a sequence labelling algorithm coupled with pattern mining, to perform postal addresses extraction from web texts, achieving the F-score of 91%. In 2013, Schmidt et al. (Schmidt et al., 2013) suggested a hybrid approach using patterns and data from gazetteers to extract addresses from web documents, showing high recall and precision. However, in their approach HTML tags are removed from the documents, which is not ideal given that geospatial information can be housed inside <meta> and other tags, as well as inside JavaScript code snippets and JSON objects. In 2018, Efremova et al. (Efremova et al., 2018) introduced a geo-tagging framework that employs n-grams, sliding windows, and a Support Vector Machine classifier to extract addresses from CC documents, achieving precision and recall of 91% and 93% respectively. Named-entity recognition tools such as spaCy191919https://spacy.io/ can also be trained to extract addresses, yet this would require a labelled training set.
Inspired by the needle-in-a-haystack evaluation task for LLMs202020This popular GitHub repository contains needle-in-a-haystack benchmark tests for LLMs: https://github.com/gkamradt/LLMTest_NeedleInAHaystack, in which long-context window LLMs are tasked to retrieve particular bits of information from a long document, such as an article or a book, we decided to employ a language model to retrieve both coordinates and addresses from web documents (if they exist), and then manually review the results. LLMs are excellent zero-shot learners and have demonstrated superior performance in named entity recognition and text classification tasks (Xie et al., 2023; Zhou et al., 2024; Wei et al., 2022; Team, 2024), so we believe they are better suited to identify the presence of geospatial in multilingual CC documents than traditional, language-and country-specific algorithms.
@@
Gemini 1.5 is one of the few LLMs currently capable of handling inputs of up to 1 million tokens (sub-word units), and has shown to perform well in the needle-in-a-haystack retrieval task (Team, 2024). The context window consideration is important given the web documents found in CC can be lengthy. Figure 1 shows the histogram of sampled document lengths (in tokens) for the documents used in our analysis. While an average document in CC, depending on the release, is between 20-50k tokens, the length distribution is heavily skewed to the right, with some web documents being as long as 1M tokens.
@@
In this paper, we use Gemini 1.5 Flash, which is a lighter and faster version of Gemini 1.5 that also has the most generous free tier API usage allowance of 1,500 requests per day. At the time of running our experiment, the pay-as-you-go tier with up to 10,000 requests per day has been announced, but is not yet available212121According to https://ai.google.dev/pricing, the pay-as-you-go pricing for Gemini 1.5 Flash and Pro will become available on May 30, 2024.
After some instruction-tuning on sampled web documents from CC, we adopted the following query:
@@
The examples of Gemini responses for queries with documents containing (1) both an address and coordinates, (2) coordinates only, and (3) address only are shown below:
@@
The document contains geographic coordinates: * **Latitude:** -7.3860078 * **Longitude:** 34.9562661 These coordinates are found within the ‘¡div class=”single_job_listing”¿‘ element.
“‘ SCIENCE MUSEUM EXHIBITION ROAD SOUTH KENSINGTON LONDON SW7 2DD “‘ “‘ https://www.google.co.uk/maps/dir//Science+Museum, +Exhibition+Rd,+Kensington,+London+SW7+2DD “‘
@@ 3.5. Determining sample size
However, in our task we are unable to utilise fast algorithms such as keyword lookup, n-grams, or K-Means. We are limited by the number of API requests we can make; in addition, each query takes Gemini between several seconds to several minutes to process. To choose the smallest yet scientifically sound sample size, we settled on Cochran’s formula (1) (Bartlett et al., 2001). The formula is used to calculate the ideal sample size that estimates population characteristics with a desired level of precision and a particular confidence interval. It is commonly used in social sciences to estimate the number of people to survey. We use the version of Cochran’s formula modified for very large (or infinite) population because a single CC release contains billions of web pages:
@@
3.6. Sampled documents
To check that our sampling is representative of the studied releases, we compare the frequency of primary document languages between our sample and the releases as reported by Common Crawl (Crawl, 2024b). The primarily language is determined by CLD2222222https://github.com/CLD2Owners/cld2.
Language frequency distribution
The prevalence of geospatial data in CC documents
English is the most common primary language of web documents in CC, representing 45.0% of the sampled releases (Crawl, 2024b) and 45.6% of our sample. The frequency of the other most common 19 languages is shown in Figure 2. We can see that our sample is balanced in relation to most common CC languages. It is worth noting that the frequency does not reflect the real-world popularity of languages (Paolillo et al., 2005). This is partly explained by the fact that the Internet originated and was popular in the Anglo-sphere for longer than the rest of the world, and because CC crawling infrastructure is located in the United States, which may force websites to serve English-language pages by default (Baack, 2024a).
| TLD | Country | Frequency (%) |
| .com | — | 45.4 |
| .org | — | 6.0 |
| .ru | Russia | 4.6 |
| .de | Germany | 3.9 |
| .net | — | 3.7 |
| .uk | United Kingdom | 2.3 |
| .jp | Japan | 2.1 |
| .fr | France | 1.6 |
| .it | Italy | 1.6 |
| .nl | Netherlands | 1.5 |
| .pl | Poland | 1.4 |
| .br | Brazil | 1.2 |
| .cn | China | 1.1 |
| .au | Australia | 1.1 |
| .edu | — | 1.1 |
Table 2 lists the most frequent top-level domains (TLDs) of the sampled documents. These are largely consistent with the summary provided by CC (Crawl, 2024b) for full releases, and the Internet-wide domain survey by W3Techs (W3Techs, 2024). The single most frequent TLD is .com, which accounts for over 45% of all sampled documents. While .com is a generic TLD that can be used by any person, organisation, or business anywhere in the world, it is especially popular in the United States in place of the country-specific .us domain.
3.7. Reviewing Gemini responses
We manually reviewed each Gemini response that is longer than 1 or 2 tokens (for ‘None’ and ‘None.’, respectively) in order to check that we are satisfied with the quality of the examples returned, as well as to distinguish whether the document contains an address, or coordinates, or both.
@@ 3.7. Reviewing Gemini responses
In order to check for false negatives, that is, the documents for which Gemini returned None that may contain geospatial data, we manually reviewed a random sample of 100 such documents. Using Google Translate232323https://translate.google.com when needed, we were unable to identify street addresses or coordinates in all cases.